A spring fastened firmly between two walls 24cm apart, find the wavelength of the second overtone.


1
Expert's answer
2021-08-02T00:19:04-0400

The possible wavelengths of the standing waves in a fixed string are given as follows (see https://brilliant.org/wiki/applying-boundary-conditions-to-standing-waves/):


"\lambda_n = \dfrac{2L}{n}"

where "L = 24cm = 0.24m" is the length of the string, and "n = 1, 2, 3..." are integer numbers. For the second overtone have:


"\lambda_2 = \dfrac{2\cdot 0.24m}{2} = 0.24m"

Answer. 0.24m.